a) The rate constant for Cr decay is 2.89×10^-7 s-¹. Calculate the time required for 87.5% decay. b) Explain the method of determining the age of an organic material. c) Explain the following terms: i) Moderator and ii) Breeder reactor.
1
Expert's answer
2020-02-12T05:48:28-0500

a)


"N(t) = N_0e^{-\lambda t}"

"875 = 1000e^{-2.89*10^{-7} t}"

"8.75 = 10e^{-2.89*10^{-7} t}"

"-2.89*10^{-7}t = ln0.875"

"-2.89*10^{-7}t = \u2212 0.13"

"t = -0.13\/(-2.89*10^{-7}) = 0.045*10^7 = 4.5*10^5 s"

b) Carbon dating - measuring the content of the radioactive isotope 14C in the material.

c)

i) Moderator - the process of reducing the kinetic energy of free neutrons as a result of their multiple collisions with the atomic nuclei of matter.

ii) Breeder reactor - nuclear reactor, allowing to produce nuclear fuel in quantities exceeding the needs of the reactor itself.